The SkyDome in Toronto, Canada has one of the world’s largest retractable roof. Its area is

32, 000 m2

, and it takes 20 min for the roof to fully retract. If you have a coil wired through the

roof, how many turns do you need to wind up your coil for you to generate an induced emf of

120 V? Assume that a uniform magnetic field with a strength of 3.0 x 10-2 T passes through

the coil parallel to its axis.


1
Expert's answer
2021-07-02T09:38:34-0400

"\xi=-n\frac Фt=Sn\frac Bt,\implies"

"n=\frac{\xi t}{SB},"

"n=\frac{120\cdot 20\cdot 60}{32000\cdot 0.03}=150."